A trolley, while going down an inclined plane, has an acceleration of 2 cm s-2. What will be its velocity 3 s after the start?
 
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


nitial Velocity of the trolley, u= 0 cms^-1
Acceleration, a= 2 cm s^-2
Time, t= 3 s
It is known that final velocity, v= u + at = 0 + 2*3 cms^-1
Therefore,
The velocity of train after 3 seconds is 6 cms^-1



Thanks